OpenStudy (anonymous):

A park in a subdivision is triangular shaped. Two adjacent sides of the park are 533 feet and 525 feet. The angle between the sides is 53°. To the nearest unit, find the area of the park in square yards

OpenStudy (anonymous):

Answer Choices: 27,935 yd^2 24,831 yd^2 37,246 yd^2 12,415 yd^2

OpenStudy (anonymous):

\[\text{area}=\frac{\text{base} \text{ * height}}{2} \]\[\text{area}\text{ = }\frac{\frac{1}{2 }525* 533 \text{ Sin}[53{}^{\circ}]}{3}\text{ = }37246.4 \text{ yd}^2 \]